# FedEx: Expediting Federated Learning over Heterogeneous Mobile Devices by Overlapping and Participant Selection

### v1

@@ 3.4. FedEx Overlapping Trigger
To avoid model drift and save the valuable resources (e.g., computing, memory and battery) of mobile devices at early training stages (Sec. 2.4), we propose a simple yet effective metric to trigger overlapping in FedEx. As discussed in Sec. 2.4, local models trained on heterogeneous data may exhibit significant differences at early FL training stages, leading to model drift, and too early overlapping/continuous computing makes it even worse, which wastes mobile devices’ resources in terms of computation, memory, and battery. Thus, it is crucial to determine when to trigger the overlapping. Here, we propose to evaluate the level of similarity between local models and the global model, and activate the overlapping when the averaged similarity scores pass a predetermined accuracy threshold. The overlapping trigger is defined as follows.
(4) $\small\frac{1}{N}\sum_{n=1}^{N}CKA(w_{n}^{r-1},w_{g}^{r-1})>\delta_{CKA},$
@@
3.5. Generalization of Staleness Ceiling
@@
As illustrated in Sec. 3.2, FedEx [^5] It is noteworthy that DGA can be regarded as a special case of FedEx with $S_{n}^{r}\leq K$, where all participating devices are homogeneous. allows device $n$ to perform at most $K$ more continuous computing iterations, i.e., $S_{n}^{r}\leq K$ in the current FL round $r$, and store a copy of the computed local model to correct for the next round. To better understand the impact of FedEx’s staleness ceiling, in this subsection, we generalize the overlapping staleness ceiling from $S_{n}^{r}\leq K$ to $S_{n}^{r}=a_{n}^{r}*K+b_{n}^{r}$, where $a_{n}^{r}$ and $b_{n}^{r}$ are two non-negative integer parameters. Based on the generalized staleness ceiling, device $n$ has to store $a_{n}^{r}$ copies of the continuous-computed local models, correct them and upload them sequentially in the next $a_{n}^{r}$ FL training rounds. Denote the upper bound of $S_{n}^{r}$ by $U$, where $U\leq\lceil(T_{n}^{cm}+T_{n,r}^{wait})/T_{n}^{cp}\rceil$.
@@
Note that $S_{n}^{r}$ plays a crucial role in balancing the pros (i.e., FL training speedup) and cons (i.e., model drift, which slows down FL convergence.) of overlapping/continuous computing. Specifically, if $S_{n}^{r}$ is smaller, the continuous computing phase is shorter, resulting in fewer local gradient descents within a certain period. This leads to a smaller acceleration effect from overlapping but also mitigates model drift. Conversely, if $S_{n}^{r}$ is larger, the continuous computing leads to larger acceleration benefits while incurring bigger model drift. Therefore, $S_{n}^{r}$ should tradeoff model drift and overlapping acceleration in order to reduce the FL training latency.
@@
From the perspective of memory consumption, as $S_{n}^{r}$ increases, $a_{n}^{r}$ monotonically increases, so that mobile device $n$ has to store more copies of staled local models with additional memory cost. That may result in excessive memory consumption, or even memory overflow, especially when device $n$ is much faster than the straggler device. By contrast, when $S_{n}^{r}$ is set to be less than certain threshold, say $K$, it ensures that only limited copies (one for $S_{n}^{r}=K$) of the local models are stored, which potentially avoids excessive memory consumption issue. Experimental results will also be presented in Sec. 5.2 to corroborate our analysis.
4. Experimental Setup
4.1. FedEx Implementation
The FedEx system has been implemented on a testbed, featuring an FL aggregator and a set of heterogeneous mobile devices. The FL server utilizes a NVIDIA RTX 3090. On the FL client side, we have incorporated 5 types of mobile devices: (1) Xiaomi 12S smartphone with Qualcomm Snapdragon 8+ Gen 1 CPU, Adreno 740 GPU, and 8GB RAM; (2) Honor 70 smartphone with Qualcomm Snapdragon 778G Plus CPU, Qualcomm Adreno 642L GPU, and 8GB RAM; (3) Honor Play 6T smartphone with MediaTed Dimensity 700 CPU, Mali-G57 MC2 GPU, and 8GB RAM; (4) NVIDIA Jetson Xavier NX with 6-core NVIDIA Carmel ARM CPU, 384-core NVIDIA Volta GPU, and 8GB RAM; (5) NVIDIA Jetson TX2 with Dual-core NVIDIA Denver CPU, 256-core NVIDIA Pascal GPU, and 4GB RAM. The FedEx system encompasses a total of 100 mobile devices, with 20 devices from each type mentioned above. For each training round, 20 devices are selected.
@@ 4.1. FedEx Implementation
FedEx is implemented by building on top of FLOWER (3). In order to realize FedEx, we need to obtain relevant parameters, such as transmission rate and training latency. To estimate on-device transmission rates, we employed the Network Monitor toolbox in the Android kernel on the smartphone, and self-developed rate monitoring module for NVIDIA devices. To measure the latency, the time of model training and transmission are recorded on mobile devices.
@@
Figure 4. Performance comparison in terms of testing accuracy and training latency under different learning tasks.
4.2. Models, Datasets and Parameters
We evaluate FedEx’s performance for three different FL tasks: image classification, next word prediction, and human activity recognition. We use four DNN models: a 2-layer CNN (26), SqueezeNet (15), MobileFormer (8), and LSTM (14), where MobileFormer is an efficient model that bridges MobileNet and Transformer.
As for the image classification task, we train a 2-layer CNN on MNIST dataset (9), a SqueezeNet on CIFAR10 dataset (17) and a MobileFormer on TSRD dataset (20). The MNIST dataset comprises 10 categories, ranging from digit “0” to “9”, and encompasses a total of 60,000 training images and 10,000 validation images. The CIFAR10 dataset comprises 50,000 training images and 10,000 validation images, distributed across 10 categories. The TSRD dataset contains 6164 traffic sign images, representing 58 sign categories. As for data distribution among mobile devices, we denote $\lambda$ as the non-i.i.d. levels, where $\lambda=0$ indicates that the data among mobile devices is i.i.d., $\lambda=0.5$ indicates that 50% of the data belong to one label and the remaining belong to other labels, and $\lambda=1$ indicates that each mobile device owns a disjoint subset of data with one label. As for the next word prediction task, we train an LSTM on the Shakespeare dataset (5), which consists of 1,129 roles and each of which is viewed as a device. Since the number of lines and speaking habits of each role vary greatly, the dataset is non-i.i.d. and unbalanced. As for the human activity recognition task, we train a 2-layer CNN on the HAR dataset (27), which is generated by having volunteers wear Samsung Galaxy S2 smartphones equipped with accelerometer and gyroscope sensors. The HAR dataset includes 10,299 data samples consisting of six categories of activities, i.e., walking, walking upstairs, walking downstairs, sitting, standing, and laying. The dataset is non-i.i.d. due to the different behavioral habits of the volunteers.
@@ 4.3. Baselines for Comparison
FedAvg (23): The FL server selects participating devices randomly, and the selected devices perform a fixed number of local training iterations.
Oort (18): The FL server selects participating devices based on the Oort utility function as defined in Eqn. (2), and the selected devices perform a fixed number of local iterations.
DGA (31): The FL server selects all mobile devices to participate in each round. The selected devices employ an overlapping training mechanism where local training continues while transmitting the model updates (the transmission starts after a fixed number of local training iterations [^6] Note that Oort/DGA is the well-recognized and top design, if not the best, in FL participant selection/overlapping FL domain, respectively. Either of them outperforms several baselines as shown in their own papers (18; 31). If the proposed FedEx performs better than DGA/Oort, it indicates that FedEx also surpasses those baselines in Oort/DGA paper..
DGAplus: The FL server selects participating devices randomly. The selected devices employ DGA policy with staleness ceiling $K$ as proposed in Sec. 3.2.
DGAplus-Oort: The FL server selects participating devices based on the Oort utility function as defined in Eqn. (2). The selected devices employ DGA policy with staleness ceiling as proposed in Sec. 3.2.
5. Evaluation Results & Analysis
5.1. Advantages in FL Training Acceleration
Table 2. Overall result of FL training with non-i.i.d. data to reach the target testing accuracy.
@@ 5.1. Advantages in FL Training Acceleration
As the evaluation results shown in Table 2 and Fig. 4, the proposed FedEx consistently outperforms its peer designs across various FL tasks, achieving remarkable training speedup. Compared with the baseline, FedAvg, FedEx expedites the FL training to the target test accuracy by approximately 1.8x, 3.0x, 2.0x, 2.1x, and 2.0x for learning tasks including CNN@MNIST, SqueezeNet@CIFAR10, MobileFormer@TSRD, LSTM@Shakespeare, and CNN@HAR, respectively. Notably, FedEx exhibits superior performance to DGA in terms of both testing accuracy and delay efficiency. As discussed in Sec. 2.2, DGA faces significant model staleness, model drift and straggler issues in heterogeneous environments, which impede or prevent the convergence of FL training. Thus, as shown in Table 2, DGA only achieves the maximum accuracy rates of 31%, 27%, 13%, 36% and 50% for the learning tasks mentioned above. In contrast, FedEx effectively manages overlap staleness and ensures timely global gradient updates. By strategically selecting appropriate devices for participation according to its overlapping aware utility function, FedEx greatly relieves the straggler effects, and achieves much better learning performance with less latency than DGA. While Oort considers both statistical utility (i.e., contributions to FL convergence) and global system utility (i.e., a latency threshold) for PS, the selected fast computing devices have to wait idly until the selected slowest device [^7] It could be one of the fast computing devices, if that mobile device has very low wireless transmission rate. finishes its current round training (computing + communications) before resuming local training for the next round. In contrast, FedEx allows for overlapping and continuous local computing, which improves the utilization of otherwise wasted waiting time, and thus further accelerates FL training. Consequently, compared with Oort, FedEx achieves a notable speedup of 1.2x, 2.0x, 1.3x, 1.3x, and 1.4x for the tested five learning tasks, respectively.
@@
Figure 5. Staleness/memory vs FL training time.
@@
Figure 6. Performance of overlapping aware PS.
5.2. Advantages of the Staleness Ceiling
@@
Table 3. Advantages of overlapping aware PS (LSTM@Shakespeare).
@@ Methods
Figure 7. FedEx performance comparison with different staleness ceilings.
Table 4. Speedup with different trigger thresholds.
@@
5.3. Advantages of Overlapping Aware PS
@@
Next, we conduct the ablation study of FedEx’s overlapping aware PS utility function, and compare FedEx with DGAplus (random PS), DGAplus-Oort (Oort PS) on the learning task LSTM@Shakespeare. As shown in Fig. 6, FedEx reaches the target accuracy with the smallest latency. Actually, it maintains the highest delay efficiency throughout the entire FL training process. That is because FedEx employs the proposed overlapping aware PS utility function, whose PS jointly considers statistical utility and overlapping latency utility, i.e., the latency reduced by overlapping/continuous computing. Conversely, the random PS in DGAplus considers neither statistical utility nor delay efficiency of FL training, and thus has the worst learning accuracy and latency performance. DGAplus-Oort directly adopts Oort’s PS utility function, and ignores the latency reduction benefits of overlapping. Thus, its performance is inferior to FedEx. It can be evidenced by the results in Table 3, where FedEx speeds up FL training to reach the target accuracy by around 1.6x compared with FedAvg, whereas DGAplus-Oort achieves only a $1.2$x acceleration over FedAvg. In addition, similar to DGA, DGAplus has the model drift issue due to overlapping at early FL training stages. Although random PS helps to mitigate it a little bit, the FL training latency in DGAplus is still $1.7$ times longer than that in FedAvg.
5.4. Analysis of FedEx Overlapping Trigger
@@
We also conduct the ablation study of FedEx overlapping trigger, and compare FedEx having different triggers with FedEx w/o trigger (i.e., FedEx overlapping right after the start of FL training), Oort and FedAvg on three different FL tasks: Mobileformer@TSRD, CNN@MNIST, and LSTM@Shakespeare. We choose trigger threshold $\delta_{CKA}$ = 0.2, 0.5, 0.7 and 0.9, respectively, and present the results in Table 4. Again, achieving a target accuracy, we let the FL training latency of FedAvg serve as the benchmark (i.e., 1x). We found that setting a too small or too big trigger threshold in FedEx may not have very good performance in latency reduction. If the trigger threshold is too small (e.g., $\delta_{CKA}$ = 0.2), participating mobile devices may start overlapping and continuous computing at very early FL training stages. That may result in model drifts, backfiring FL convergence speed and wasting the computing and memory resources of mobile devices, which verify our analysis in Sec. 3.4. The same reason also applies to FedEx w/o trigger. If the trigger is too big (e.g., $\delta_{CKA}$ = 0.9), it may be too late to start overlapping, so that FedEx has missed many opportunities for further latency reduction. From the results in Table 4, among the chosen FedEx trigger values, $\delta_{CKA}$ = 0.7 yields the best speedup performance, i.e., 2.0x speedup and overlapping starts at Round 207 for MobileFormer@TSRD, 1.8x speedup and overlapping starts at Round 39 for CNN@MNIST, and 2.1x speedup and overlapping starts at Round 82 for LSTM@Shakespeare. Although $\delta_{CKA}$ = 0.7 has good performance, it may not be the optimal overlapping trigger in FedEx. We will leave how to set the optimal trigger for our future studies.
5.5. Impacts of Different Staleness Ceilings

### v3

@@ III-D FedEx Overlapping Trigger
To avoid model drift and save the valuable resources (e.g., computing, memory and battery) of mobile devices at early training stages (Sec. II-D), we propose a simple yet effective metric to trigger overlapping in FedEx. As discussed in Sec. II-D, local models trained on heterogeneous data may exhibit significant differences at early FL training stages, leading to model drift, and too early overlapping/continuous computing makes it even worse, which wastes mobile devices’ resources in terms of computation, memory, and battery. Thus, it is crucial to determine when to trigger the overlapping. Here, we propose to evaluate the level of similarity between local models and the global model, and activate the overlapping when the averaged similarity scores pass a predetermined accuracy threshold. The overlapping trigger is defined as follows.
$\small\frac{1}{N}\sum_{n=1}^{N}CKA(w_{n}^{r-1},w_{g}^{r-1})>\delta_{CKA},\vspace{-2mm}$ (4)
@@
III-E Generalization of Staleness Ceiling
@@
As illustrated in Sec. III-B, FedEx [^5] It is noteworthy that DGA can be regarded as a special case of FedEx with $S_{n}^{r}\leq K$, where all participating devices are homogeneous. allows device $n$ to perform at most $K$ more continuous computing iterations, i.e., $S_{n}^{r}\leq K$ in the current FL round $r$, and store a copy of the computed local model to correct for the next round. To better understand the impact of FedEx’s staleness ceiling, in this subsection, we generalize the overlapping staleness ceiling from $S_{n}^{r}\leq K$ to $S_{n}^{r}=a_{n}^{r}*K+b_{n}^{r}$, where $a_{n}^{r}$ and $b_{n}^{r}$ are two non-negative integer parameters. Based on the generalized staleness ceiling, device $n$ has to store $a_{n}^{r}$ copies of the continuous-computed local models, correct them and upload them sequentially in the next $a_{n}^{r}$ FL training rounds. Denote the upper bound of $S_{n}^{r}$ by $U$, where $U\leq\lceil(T_{n}^{cm}+T_{n,r}^{wait})/T_{n}^{cp}\rceil$.
@@
Note that $S_{n}^{r}$ plays a crucial role in balancing the pros (i.e., FL training speedup) and cons (i.e., model drift, which slows down FL convergence) of overlapping/continuous computing. Specifically, if $S_{n}^{r}$ is smaller, the continuous computing phase is shorter, resulting in fewer local gradient descents within a certain period. This leads to a smaller acceleration effect from overlapping but also mitigates model drift. Conversely, if $S_{n}^{r}$ is larger, the continuous computing leads to larger acceleration benefits while incurring bigger model drift. Therefore, $S_{n}^{r}$ should tradeoff model drift and overlapping acceleration in order to reduce the FL training latency.
@@
From the perspective of memory consumption, as $S_{n}^{r}$ increases, $a_{n}^{r}$ monotonically increases, so that mobile device $n$ has to store more copies of staled local models with additional memory cost. That may result in excessive memory consumption, or even memory overflow, especially when device $n$ is much faster than the straggler device. By contrast, when $S_{n}^{r}$ is set to be less than certain threshold, say $K$, it ensures that only limited copies (one for $S_{n}^{r}=K$) of the local models are stored, which potentially avoids excessive memory consumption issue. Experimental results will also be presented in Sec. V-B to corroborate our analysis.
IV Experimental Setup
IV-A FedEx Implementation
@@
The FedEx system has been implemented on a testbed, featuring an FL aggregator and a set of real-world heterogeneous mobile devices, detailed in the Appendix E. FedEx is implemented by building on top of FLOWER [17]. In order to realize FedEx, we need to obtain relevant parameters, such as transmission rate and training latency. To estimate on-device transmission rates, we employed the Network Monitor toolbox in the Android kernel on the smartphone, and self-developed rate monitoring module for NVIDIA devices. To measure the latency, the time of model training and transmission are recorded on mobile devices.
@@
Figure 4: Performance comparison in terms of testing accuracy and training latency under different learning tasks.
IV-B Models, Datasets and Parameters
We evaluate FedEx’s performance for three different FL tasks: image classification, next word prediction, and human activity recognition. We use four DNN models: a 2-layer CNN [18], SqueezeNet [19], MobileFormer [20], and LSTM [21], where MobileFormer is an efficient model that bridges MobileNet and Transformer.
As for the image classification task, we train a 2-layer CNN on MNIST dataset [22], a SqueezeNet on CIFAR10 dataset [23] and a MobileFormer on TSRD dataset [24]. The MNIST dataset comprises 10 categories, ranging from digit “0” to “9”, and encompasses a total of 60,000 training images and 10,000 validation images. The CIFAR10 dataset comprises 50,000 training images and 10,000 validation images, distributed across 10 categories. The TSRD dataset contains 6164 traffic sign images, representing 58 sign categories. As for data distribution among mobile devices, we denote $\lambda$ as the non-i.i.d. levels, where $\lambda=0$ indicates that the data among mobile devices is i.i.d., $\lambda=0.5$ indicates that 50% of the data belong to one label and the remaining belong to other labels, and $\lambda=1$ indicates that each mobile device owns a disjoint subset of data with one label. As for the next word prediction task, we train an LSTM on the Shakespeare dataset [25], which consists of 1,129 roles and each of which is viewed as a device. Since the number of lines and speaking habits of each role vary greatly, the dataset is non-i.i.d. and unbalanced. As for the human activity recognition task, we train a 2-layer CNN on the HAR dataset [26], which is generated by having volunteers wear Samsung Galaxy S2 smartphones equipped with accelerometer and gyroscope sensors. The HAR dataset includes 10,299 data samples consisting of six categories of activities, i.e., walking, walking upstairs, walking downstairs, sitting, standing, and laying. The dataset is non-i.i.d. due to the different behavioral habits of the volunteers.
@@ IV-C Baselines for Comparison
FedAvg [27]: The FL server selects participating devices randomly, and the selected devices perform a fixed number of local training iterations.
Oort [15]: The FL server selects participating devices based on the Oort utility function as defined in Eqn. (2), and the selected devices perform a fixed number of local iterations.
DGA [9]: The FL server selects all mobile devices to participate in each round. The selected devices employ an overlapping training mechanism where local training continues while transmitting the model updates (the transmission starts after a fixed number of local training iterations [^6] Note that Oort/DGA is the well-recognized and top design, if not the best, in FL participant selection/overlapping FL domain, respectively. Either of them outperforms several baselines as shown in their own papers [15, 9]. If the proposed FedEx performs better than DGA/Oort, it indicates that FedEx also surpasses those baselines in Oort/DGA paper.).
DGAplus: The FL server selects participating devices randomly. The selected devices employ DGA policy (overlapping training) with staleness ceiling as proposed in Sec. III-B.
DGAplus-Oort: The FL server selects participating devices based on the Oort utility function as defined in Eqn. (2). The selected devices employ DGA policy (overlapping training) with staleness ceiling as proposed in Sec. III-B.
V Evaluation Results & Analysis
V-A Advantages in FL Training Acceleration
TABLE II: Overall result of FL training with non-i.i.d. data to reach the target testing accuracy.
@@ V-A Advantages in FL Training Acceleration
As the evaluation results shown in Table II and Fig. 4, the proposed FedEx consistently outperforms its peer designs across various FL tasks, achieving remarkable training speedup. Compared with the baseline, FedAvg, FedEx expedites the FL training to the target test accuracy by approximately 1.8x, 3.0x, 2.0x, 2.1x, and 2.0x for learning tasks including CNN@MNIST, SqueezeNet@CIFAR10, MobileFormer@TSRD, LSTM@Shakespeare, and CNN@HAR, respectively. Notably, FedEx exhibits superior performance to DGA in terms of both testing accuracy and delay efficiency. As discussed in Sec. II-B, DGA faces significant model staleness, model drift and straggler issues in heterogeneous environments, which impede or prevent the convergence of FL training. Thus, as shown in Table II, DGA only achieves the maximum accuracy rates of 31%, 27%, 13%, 36% and 50% for the learning tasks mentioned above. In contrast, FedEx effectively manages overlap staleness and ensures timely global gradient updates. By strategically selecting appropriate devices for participation according to its overlapping aware utility function, FedEx greatly relieves the straggler effects, and achieves much better learning performance with less latency than DGA. While Oort considers both statistical utility (i.e., contributions to FL convergence) and global system utility (i.e., a latency threshold) for PS, the selected fast computing devices have to wait idly until the selected slowest device [^7] It could be one of the fast computing devices, if that mobile device has very low wireless transmission rate. finishes its current round training (computing + communications) before resuming local training for the next round. In contrast, FedEx allows for overlapping and continuous local computing, which improves the utilization of otherwise wasted waiting time, and thus further accelerates FL training. Consequently, compared with Oort, FedEx achieves a notable speedup of 1.2x, 2.0x, 1.3x, 1.3x, and 1.4x for the tested five learning tasks, respectively.
@@
Figure 5: Staleness/memory vs FL training time.
@@
Figure 6: Performance of overlapping aware PS.
V-B Advantages of the Staleness Ceiling
@@
TABLE III: Advantages of overlapping aware PS (LSTM@Shakespeare).
@@ Methods
Figure 7: FedEx performance comparison with different staleness ceilings.
TABLE IV: Speedup with different trigger thresholds.
@@
V-C Advantages of Overlapping Aware PS
@@
Next, we conduct the ablation study of FedEx’s overlapping aware PS utility function, and compare FedEx with DGAplus (random PS), DGAplus-Oort (Oort PS) on the learning task LSTM@Shakespeare. As shown in Fig. 6, FedEx reaches the target accuracy with the smallest latency. Actually, it maintains the highest delay efficiency throughout the entire FL training process. That is because FedEx employs the proposed overlapping aware PS utility function, whose PS jointly considers statistical utility and overlapping latency utility, i.e., the latency reduced by overlapping/continuous computing. Conversely, the random PS in DGAplus considers neither statistical utility nor delay efficiency of FL training, and thus has the worst learning accuracy and latency performance. DGAplus-Oort directly adopts Oort’s PS utility function, and ignores the latency reduction benefits of overlapping. Thus, its performance is inferior to FedEx. It can be evidenced by the results in Table III, where FedEx speeds up FL training to reach the target accuracy by around 1.6x compared with FedAvg, whereas DGAplus-Oort achieves only a $1.2$x acceleration over FedAvg. In addition, similar to DGA, DGAplus has the model drift issue due to overlapping at early FL training stages. Although random PS helps to mitigate it a little bit, the FL training latency in DGAplus is still $1.7$ times longer than that in FedAvg.
V-D Analysis of FedEx Overlapping Trigger
@@
We also conduct the ablation study of FedEx overlapping trigger, and compare FedEx having different triggers with FedEx w/o trigger (i.e., FedEx overlapping right after the start of FL training), Oort and FedAvg on three different FL tasks: Mobileformer@TSRD, CNN@MNIST, and LSTM@Shakespeare. We choose trigger threshold $\delta_{CKA}$ = 0.2, 0.5, 0.7 and 0.9, respectively, and present the results in Table IV. Again, achieving a target accuracy, we let the FL training latency of FedAvg serve as the benchmark (i.e., 1x). We found that setting a too small or too big trigger threshold in FedEx may not have very good performance in latency reduction. If the trigger threshold is too small (e.g., $\delta_{CKA}$ = 0.2), participating mobile devices may start overlapping and continuous computing at very early FL training stages. That may result in model drifts, backfiring FL convergence speed and wasting the computing and memory resources of mobile devices, which verify our analysis in Sec. III-D. The same reason also applies to FedEx w/o trigger. If the trigger is too big (e.g., $\delta_{CKA}$ = 0.9), it may be too late to start overlapping, so that FedEx has missed many opportunities for further latency reduction. From the results in Table IV, among the chosen FedEx trigger values, $\delta_{CKA}$ = 0.7 yields the best speedup performance, i.e., 2.0x speedup and overlapping starts at Round 207 for MobileFormer@TSRD, 1.8x speedup and overlapping starts at Round 39 for CNN@MNIST, and 2.1x speedup and overlapping starts at Round 82 for LSTM@Shakespeare. Although $\delta_{CKA}$ = 0.7 has good performance, it may not be the optimal overlapping trigger in FedEx. We will leave how to set the optimal trigger for our future studies.
V-E Impacts of Different Staleness Ceilings